### v1

# Advances in Automated Fetal Brain MRI Segmentation and Biometry: Insights from the FeTA 2024 Challenge

Abstract
Accurate segmentation and biometric analysis are essential for studying the developing fetal brain in utero. The Fetal Brain Tissue Annotation (FeTA) Challenge 2024 builds upon previous editions to further advance the clinical relevance and robustness of automated fetal brain MRI analysis. This year’s challenge introduced biometry prediction as a new task complementing the usual segmentation task. The segmentation task also included a new low-field (0.55T) MRI testing set and used Euler characteristic difference (ED) as a topology-aware metric for ranking, extending the traditional overlap or distance-based measures.
A total of 16 teams submitted segmentation methods for evaluation. Segmentation performance across top teams was highly consistent across both standard and low-field MRI data. Longitudinal analysis over past FeTA editions revealed minimal improvement in accuracy over time, suggesting a potential performance plateau, particularly as results now approach or surpass reported levels of inter-rater variability. However, the introduction of the ED metric revealed topological differences that were not captured by conventional metrics, underscoring its value in assessing segmentation quality. Notably, the curated low-field MRI dataset achieved the highest segmentation performance, illustrating the potential of affordable imaging systems when combined with high-quality preprocessing and reconstruction.
A total of 7 teams submitted automated biometry methods for evaluation. While promising, this task exposed a critical limitation: most submitted methods failed to outperform a simple baseline that predicted measurements based solely on gestational age, without using image data. Performance varied widely across biometric measurements and between teams, indicating both current challenges and opportunities for improvement in this area. These findings highlight the need for better integration of volumetric context and stronger modeling strategies needed for the clinical adoption of automated fetal biometry estimation.
In addition, we analyzed different dimensions of domain shifts within our data and observed that image quality was the most influential factor affecting model generalization, with Dice score differences of up to 0.10 between low- and high-quality scans. The choice of super-resolution reconstruction pipeline also had a substantial impact on segmentation performance. Other factors—such as gestational age, pathology, and acquisition site—also contributed to performance variability, but their effects were comparatively smaller.
Overall, FeTA 2024 provides a rigorous, multi-faceted benchmark for evaluating multi-class segmentation and biometry estimation in fetal brain MRI. It emphasizes the need for data-centric approaches, improved topological modeling, and greater dataset diversity to develop clinically reliable and generalizable AI tools for fetal neuroimaging.
keywords:
Fetal Brain, MRI, Low-field, Segmentation, Topology, Biometry, Domain Shift, Challenge results
journal: Medical Image Analysis
[…]
[40]organization=Pattern Analysis and Learning Group, Department of Radiation Oncology, Heidelberg University Hospital, city=Heidelberg, country=Germany
[…]
[41]organization=Helmholtz Imaging, German Cancer Research Center (DKFZ), city=Heidelberg, country=Germany
[…]
[44]organization=Croatian Institute for Brain Research, School of Medicine, University of Zagreb, city=Zagreb, country=Croatia
[…]
Equal contributions.
1 Introduction
The fetal brain undergoes rapid and complex development throughout gestation, influenced by both genetic and environmental factors. Understanding this dynamic process is critical in both clinical and research domains, as neurodevelopmental disruptions are linked to congenital anomalies and long-term cognitive or physiological impairments (Griffiths et al., 2017; Ciceri et al., 2024; Van den Bergh et al., 2018). In vivo imaging biomarkers derived from ultrasonography (US) or magnetic resonance imaging (MRI) provide non-invasive and quantifiable metrics to monitor prenatal brain development. Deviations from normative patterns in these biomarkers have been associated with a range of pathologies, including corpus callosum (Marathu et al., 2024; Lamon et al., 2024) and posterior fossa malformations (Dovjak et al., 2020; Mahalingam et al., 2021), ventriculomegaly (Chen et al., 2024), and have been shown to correlate with neurodevelopmental outcomes in conditions such as congenital heart disease (Sadhwani et al., 2022), intrauterine growth restriction (Egaña-Ugrinovic et al., 2015; Meijerink et al., 2023), and preterm birth (Story et al., 2021; Hall et al., 2024).
Fetal brain MRI has emerged as an important non-invasive tool for studying neurodevelopment in utero and diagnosing congenital disorders, complementing ultrasonography (Griffiths et al., 2017; Alamo et al., 2010). Accurate and automatic segmentation of fetal brain tissues in MRI is critical for quantitative analysis and biomarker extraction, including tissue volumetry, cortical morphometry (Payette et al., 2023), and biometric measurements (She et al., 2023). Manual segmentation, however, remains labor-intensive, error-prone, and susceptible to inter-observer variability, underscoring the necessity of reliable automated techniques.
While clinical US and 2D MRI are the standard techniques for assessing fetal development (Tilea et al., 2009), the use of super-resolution reconstruction (SRR) techniques to generate 3D fetal brain reconstructions has emerged as a powerful advancement. SRR methods fuse multiple 2D MRI slices (often motion-corrupted) into a single, enhanced 3D motion-corrected volume, significantly improving brain analysis (Gafner et al., 2020; Avisdris et al., 2021; Matthew et al., 2024). Recent studies have shown that biometric measurements derived from 3D SRR volumes correlate strongly with those from ultrasound, while offering greater rater confidence than using 2D MRI series (Lamon et al., 2024; Khawam et al., 2021; Gafner et al., 2020; Sanchez et al., 2024b; Kyriakopoulou et al., 2016; Ciceri et al., 2023).
The Fetal Tissue Annotation (FeTA) challenges, held in 2021 (Payette et al., 2023) and 2022 (Payette et al., 2024), have significantly advanced fetal brain MRI analysis by providing public datasets and standardized evaluation protocols for brain tissue segmentation. The FeTA 2024 challenge builds on previous editions, retaining the core brain tissue segmentation task and introducing a new clinically relevant objective: biometry extraction, alongside several other key innovations.
Firstly, FeTA 2024 introduces a new low-field (LF, 0.55T) MRI testing dataset. LF MRI offers a low-cost alternative to 1.5–3T systems, making it especially valuable in resource-limited settings (Arnold et al., 2023; Marques et al., 2019). This affordability supports research in low- and middle-income countries with large pediatric populations, where access to high-field MRI is limited, hindering studies on brain development under normal and adverse conditions (Murali et al., 2023; Aviles Verdera et al., 2023).
Secondly, we introduced the Euler characteristic difference as an additional ranking metric for segmentation (Taha and Hanbury, 2015). Unlike overlap- or distance-based metrics, it captures topological correctness, offering a complementary view of performance (Maier-Hein et al., 2024). This is especially relevant for downstream tasks like cortical surface extraction or morphometric analyses (e.g., sulcal folding, cortical maturation, or structural abnormality assessment) (Yehuda et al., 2023; Clouchoux et al., 2011).
In FeTA 2024, we promote the development of generalizable, fully automated methods for fetal brain analysis, enabling the extraction of key imaging biomarkers via multi-class tissue segmentation and biometry across diverse acquisition and reconstruction settings. This paper offers a comprehensive overview of the challenge, covering its organization, submitted algorithms, performance assessment, benchmarking, and evaluation using the BIAS reporting framework, which emphasizes transparency, reproducibility, and fairness (Maier-Hein et al., 2020). We also analyze performance trends over time, tracking improvements in state-of-the-art segmentation accuracy across FeTA editions. Finally, we assess data quality in both training and testing sets to examine its impact on the generalization of submitted methods. Combined with other domain shifts—such as gestational age, pathology, super-resolution reconstruction, and acquisition site—our work provides a deep overview of how domain shifts affect deep learning models for fetal brain analysis and informs strategies to mitigate their impact.
2 Methods
2.1 Challenge organization
Context
The FeTA 2024 challenge was held as a thematic event within the Perinatal, Preterm, and Pediatric Image Analysis (PIPPI) workshop [^1] https://pippiworkshop.github.io/, part of the Medical Image Computing and Computer-Assisted Intervention (MICCAI) 2024 conference. The challenge was run through a custom platform, available at https://fetachallenge.github.io/, which provided participants with all the necessary information on the organization, time frame, and submission instructions.
Data, participation and submission
Challenge participation required submission of fully automated segmentation and/or biometry algorithms. A training set of 3D super-resolution fetal brain MRI from two institutions was provided; no validation set was released, and test data remained private for evaluation. Participants could use publicly available external datasets and pre-trained models, provided these were public and fully documented in the algorithm description, as well as use both 2D and 3D models.
Participants submitted their algorithms as Docker containers with a command-line interface for test data evaluation [^2] Instructions: https://github.com/fetachallenge/fetachallengesubmission.. Any programming language was allowed, provided the input/output followed the evaluation utility specifications. Each team was allowed one submission, except in cases of technical errors (e.g., Docker issues), which could be corrected upon notification by the organizers. Evaluation on test data was performed by the organizers using publicly available code [^3] Available at https://fetachallenge.github.io/pages/Evaluation.. To promote transparency and reproducibility, FeTA 2024 encouraged participants to share their code publicly. A Docker Hub page (https://hub.docker.com/repositories/fetachallenge2024) was created to host containers from teams who agreed to release their Docker images.
Timetable, rewards and results paper
The challenge followed a predefined schedule: training data was released on May 21, 2024; registration opened after challenge acceptance. The Docker submission deadline was extended to August 4, 2024, and algorithm descriptions were due by August 12. On August 23, the top five teams were invited to prepare 2-minute pitch presentations for the challenge day and, along with all participants, were invited to present posters at the dedicated conference session. The challenge took place in person on October 6, 2024, during MICCAI. Results were announced live and later published on the challenge website, along with top teams’ presentations (with their consent). The top three teams in each task received certificates and small gifts, including a 3D-printed fetal brain keychain for in-person attendees. The highest-ranking team in each task also received a box of artisanal Swiss-made cookies. Organizers could participate but were not eligible for awards.
All teams with valid submissions and interest in the publication were included in this results paper, with up to three members per team listed as co-authors. Teams were free to publish their algorithms and results independently after the challenge, without embargo, provided they cited both the data publication (Payette et al., 2021) and this summary paper.
Data usage terms and conflicts of interest
The training data from the University Children’s Hospital Zürich (Kispi) and General Hospital Vienna/Medical University of Vienna were provided with specific licensing conditions. Kispi data, hosted on the Synapse platform [^4] https://www.synapse.org/Synapse:syn25649159/wiki/610007, is for non-commercial use only. Vienna data is governed by a custom Data Transfer Agreement, allowing use for challenge purposes only. Participants could modify the data, including generating synthetic data through augmentation, as long as modifications were documented and synthetic data could be provided to the organizers upon request. None of the organizers participated in this year’s challenge or have conflicts of interest to disclose. The challenge awards were funded by the institutional budget (Kispi), and none of the participants were involved in funding. Only organizers at Kispi had full access to the testing dataset, as they managed data transfer agreements with all providers.
2.2 Challenge tasks
The FeTA challenge presents two primary tasks (see Figure 1). Participants could choose to compete in either or both tasks.
Task 1. Fetal brain tissue segmentation
This task aims to develop algorithms that automatically delineate different tissues in SRR fetal brain MRI. The 3D semantic segmentation involves classifying each voxel into one of seven predefined classes: Background, External CSF, Grey Matter (GM), White Matter (WM), Ventricles including cavum (VM), Cerebellum (CBM), Deep Grey Matter (SGM), and Brainstem (BSM). Reference annotation procedures and inter-rater variability analyses for all datasets (except the new LF set) are detailed in Payette et al. (2021) and Payette et al. (2024). The LF dataset followed the same annotation protocol, with seven annotators (AJ, CS, RG, VZ, YG, MA, MR) each segmenting a specific label map. These were merged into a single reference annotation, reviewed, and corrected by two fetal MRI experts (KP, AJ).
Task 2. Biometric measurements prediction
The goal of this task is to develop algorithms that automatically and accurately estimate key fetal brain biometry from MRI. The selected measurements—length of the corpus callosum (LCC), height of the vermis (HV), brain biparietal diameter (bBIP), skull biparietal diameter (sBIP), and transverse cerebellar diameter (TCD)—were chosen to minimize annotation burden while providing complementary anatomical and diagnostic value. Four raters contributed: YG (5 years’ fetal MRI experience), MKo (16 years), and junior raters RG and MA (reviewed by AJ, 12 years) [^5] Biometry annotation protocol is described on our website: https://fetachallenge.github.io/pages/Data_description. Not all measurements were available for all cases: in the test set, 15 cases lacked LCC, one lacked HV, and one lacked TCD due to annotator uncertainty. In the training set, 102 of 120 cases had complete annotations—10 Kispi cases were excluded for poor quality; 5 Kispi and 3 Vienna cases had partial annotations. While the main goal was to predict biometry values, the training set also included 3D landmark annotations—single-voxel labels marking anatomical structures used to derive each measurement. Clinicians identified these landmarks during annotation, and the actual biometry values were computed via organizer-provided scripts. Both the landmarks and scripts were shared, allowing participants to either regress biometry directly or predict landmarks, followed by automated biometry measurement.
Figure 1: FeTA 2024 Challenge Tasks. Task 1 involves segmentation of fetal brain tissues into seven classes, while Task 2 focuses on estimating five biometric measurements, both illustrated in the figure.
2.3 Challenge data sets
Subject selection aimed to ensure a representative cohort spanning 18–35 weeks of gestation, including both neurotypical and pathological cases (e.g., spina bifida, ventriculomegaly, corpus callosum malformations) to reflect clinical practice. UCSF and CHUV data were acquired during clinical fetal MRI scans following ultrasound referral, performed by trained medical staff. Data from KCL, Kispi, and Vienna were collected using research protocols. All cohorts had approval by the local ethics committee for use in the challenge after anonymization [^6] KISPI: Ethical Committee of the Canton of Zurich, Switzerland (Decision numbers: 2017 00885, 2016 01019, 2017 00167). CHUV: Ethics Committee of the Canton de Vaud, Switzerland (CER-VD 2021 00124). Vienna: Approved by the ethics review board and data clearing department at the Medical University of Vienna. UCSF: institutional review board (IRB 16 20619). KCL: Ethics Committee Dulwich (Ethics code 19 LO 0852)..
Each case included a 3D fetal brain MRI reconstruction, manual brain tissue segmentation, and biometry annotations. Metadata included gestational age (GA) and a binary label indicating neurotypical or pathological status. To preserve anonymity, gender was excluded and GA was randomly offset by ±3 days.
Figure 2: FeTA 2024 data distribution by GA (weeks), condition, and image quality (0 = lowest, 4 = highest; 1 = minimum acceptable), stratified by Site and SR method for training (top) and testing (bottom) sets. In the image quality plots, the red dotted line marks the threshold score (1.0); images with a score below this value are classified as poor quality.
The challenge dataset comprises of 120 training and 180 test cases. The test set was split into in-domain (from the same institutions and protocols as training data) and out-of-domain cases. To ensure balance, both subsets were similar in size to the training set. Demographic characteristics, including GA and pathology distribution, were matched across training and testing cohorts (see Figure 2). For this year’s challenge, manual quality control was performed on all training and testing cases following the protocol by Sanchez et al. (2024a), ensuring comparable data quality between training and testing sets (Figure 2).
The FeTA 2024 training and testing datasets are identical to those used in FeTA 2022, with two additions: a new low-field out-of-domain test set from King’s College London (KCL) and manual biometry annotations.
All cases were acquired using T2-weighted single-shot fast spin-echo sequences [^7] Also known as HASTE (Siemens), SSTSE (Philips), or SSFSE (GE), depending on the scanner manufacturer., the standard for structural fetal MRI due to their high signal-to-noise ratio and reduced sensitivity to fetal motion. To further mitigate motion artifacts, multiple stacks were acquired in various orientations (axial, sagittal, coronal, and off-plane). Manual selection of 2D stacks was done at each site and then combined into a single high-resolution, isotropic 3D image via super-resolution reconstruction. The resulting 3D volumes were zero-padded to 256x256x256 and reoriented to a standard radiological plane. A summary of acquisition parameters, demographic characteristics, and reconstruction methods in all sites is provided in Table 1. Additional details about the new KCL dataset are provided below. For further detailed information on the FeTA 2022 acquisitions, please refer to Payette et al. (2024).
Table 1: FeTA 2024 datasets properties. $N_{n}$ number of neurotypical subjects, $N_{p}$ number of pathological subjects. ”+” indicates the minimum TE value
Used Testing Institution Scanner N SR SR res. TR/TE GA $\mathbf{N_{n}/N_{p}}$
for domain method (mm3) (ms) (weeks)
Training In domain KISPI
GE Signa Discovery
MR450/MR750
(1.5T/3T respectively)*
80
MIALSRTK (40)
IRTK-simple (40)
(0.5)3
2000-3500/
120+
20-34.4 49/31
Vienna
Philips Ingenia/Intera (1.5T)
Philips Achieva (3T)*
40 NiftyMIC (1.0)3
6000-22000/
80-140
19.3-34.4 33/7
Testing In domain KISPI
GE Signa Discovery
MR450/MR750
(1.5T/3T respectively)*
40
MIALSRTK (20)
IRTK-simple (20)
(0.5)3
2000-3500/
120+
21.3-34.6 15/25
Vienna
Philips Ingenia/Intera (1.5T)
Philips Achieva (3T)*
40 NiftyMIC (1.0)3
6000-22000/
80-140
18.1-35.5 20/20
Out of domain CHUV
Siemens MAGNETOM
Aera (1.5T)
40 MIALSRTK (1.125)3 1200/90 21.0-35.0 25/15
UCSF
GE Signa Discovery
MR750/MR750W (3T)
40 NiftyMIC (0.8)3
200-3500/
100+
20.0-35.1 8/32
KCL
Siemens MAGNETOM
Free.Max (0.55T)
20 SVRTK (0.8)3 2500/106 21.0-35.0 15/5
*The training dataset contained data from both 1.5T and 3T scanners. However, which cases belonged to which scanner were not provided to the participants as it was part of the data anonymization process. Therefore, the breakdown of number of cases per scanner is not provided here.
Data from KCL was collected using a 0.55T low-field MRI scanner (Siemens MAGNETOM Free.Max) with a HASTE sequence as part of a prospective single-center study and fully anonymized following local procedures (Ethics Committee Dulwich 19 LO 0852). The acquired stacks had a resolution of 1.5mm x 1.5mm x 4.5mm, which were then reconstructed into a high-resolution volume of 0.8mm x 0.8mm x 0.8mm using SVRTK (Uus et al., 2020). Key acquisition parameters include a flip angle of 180°, a field-of-view of 450 × 450mm², and a base resolution of 304x304 pixels, yielding a voxel size of 1.5 × 1.5 × 4.5mm³. The acquisition time ranged from 64 to 122 seconds. Data collection took place at St Thomas’ Hospital in London, United Kingdom, without the use of maternal or fetal sedation. All acquisitions were performed using the contour L coil and the integrated spine coil while the mother was in a supine position. This dataset is used only in the testing set.
2.4 Evaluation Metrics
We provide a short recall of the ranking metrics. The detailed mathematical formulation is available in supplementary materials A1.
Task 1. Segmentation
Performance of segmentation algorithms is comprehensively assessed through complementary metrics of spatial overlap, volume, shape, and topological correctness:
  • 1.
    Dice Similarity Coefficient (Dice; $\uparrow$)888$\uparrow$ means that a higher score is better and $\downarrow$ that a lower score is better: measures voxel-wise correspondence between the predicted and ground truth (GT) segmentations.
  • 2.
    Volume Similarity (VS; $\uparrow$): measures the similarity of the volumes between the predicted and GT segmentations.
  • 3.
    Hausdorff Distance (HD95; $\downarrow$): quantifies the distance between contours of the predicted and GT segmentations with robustness to outliers.
  • 4.
    Euler Characteristics Difference (ED; $\downarrow$): evaluates the topological similarity between the predicted and GT segmentations.
As ED is included in the ranking for the first time, we describe it further. It is based on the Euler characteristic (EC):
$EC=\text{BN}_{0}-\text{BN}_{1}+\text{BN}_{2}$
where Betti Number $\text{BN}_{0}$ represents the number of connected components (i.e., regions), $\text{BN}_{1}$ represents the number of loops or holes and $\text{BN}_{2}$ represents the number of voids or cavities. The ED difference is then computed as $|EC_{pred}-EC_{GT}|$. Smaller differences indicate better topological alignment. The Betti number values of GT are: for all brain tissue labels, $\text{BN}_{1}=0$ and $\text{BN}_{2}=0$. For the eCSF, WM, ventricles, cerebellum, dGM, and brainstem, $\text{BN}_{0}=1$, while for GM, $\text{BN}_{0}=2$.
Task 2: Biometry Estimation
The primary metric for evaluating biometry estimation algorithms is mean average percentage error (MAPE; $\downarrow$) , which quantifies the error in the estimated biometric measurements relative to the actual measurements:
$MAPE=\frac{1}{N}\sum_{i=1}^{N}\frac{|y_{i}-\hat{y}_{i}|}{\text{$y_{i}$}}\times 100,$
where $y_{i}$ and $\hat{y}_{i}$ are GT and predicted measurements respectively, and $N$ is the total number of measurements. This metric accounts for variable sizes of the target structures and is used to assess the accuracy of the estimated biometric measurements.
2.5 Ranking
Submissions are ranked based on metrics computed for each brain tissue label (or biometric measurement) in the predicted maps of the fetal brain volumes. For segmentation, the final rank is the average of all 4 metrics: Dice, HD95, VS, and ED. For biometry, the final rank is based on MAPE. For metrics where higher values are better (Dice, VS), the algorithm with the highest value ranks best. For metrics where lower values are better (like HD95 and ED for segmentation and MAPE for biometry tasks), the algorithm with the lowest value ranks best. The individual label rankings are summed, and the algorithm with the highest combined rank is considered the best.
In cases of missing results (e.g., if an algorithm fails to detect a label or if the entire label map is empty), the worst possible values will be assigned to the algorithm. For example, if a label is missing in the label map, it will receive a Dice and VS of 0. For HD95, EC, and MAPE, the missing values are set to double the maximum value of other algorithms for that sub-ranking. This ensures that algorithms with missing results are ranked last for that specific task/brain tissue.
2.5.1 Biometry baselines
In the ranking of Task 2, two additional baseline models representing lower and upper performance limits were incorporated as separate submissions. These entries, intended solely for benchmarking purposes, were not considered in the formal determination of the challenge competition ranking.
Lower bound: Gestational age regression model
This model, referred to as [GA] in the result’s Table 5, is a simple univariate linear regression baseline. For each biometric measurement $y$, the model predicts its value $\hat{y}$ using the gestational age (GA) as the sole explanatory variable, mathematically:
$\hat{y}=\beta_{0}+\beta_{1}\cdot\text{GA},$
where $\beta_{0}$ is the intercept and $\beta_{1}$ is the regression coefficient learned from the training data. This baseline does not rely on the image and aims at quantifying how strongly the GA can account for the size of a given structure.
Upper bound: Inter-rater variability
The upper bound is set by averaging inter-rater variability, further denoted as [inter-rater]. This reflects the best-expected accuracy, accounting for measurement errors and uncertainties between manual raters. For each biometric measurement in the test dataset, annotations from two independent observers are used by comparing one observer’s measurement to the other’s, with the result averaged across all test cases.
2.6 Statistical analysis
The non-parametric Wilcoxon signed-rank test was used to assess performance differences between algorithms, as the Shapiro-Wilk test indicated non-normal distribution. To evaluate performance differences across subsets (e.g., neurotypical vs. pathological cohorts), we applied the Mann-Whitney U test (Wilcoxon rank-sum test). For all tests, statistical significance was set at $p<0.05$. For multiple comparisons, such as between sites or labels, we applied Bonferroni correction.
2.7 Further analysis
FeTA 2024, as the third edition of the challenge, provides an opportunity to assess progress and unsolved challenges. We report two additional analyses: (i) the evolution of top-performing segmentation models over the last three editions, and (ii) the impact of different domain shift sources on model performance.
2.7.1 Insights from three years of competition: progress or plateau?
To assess progress in fetal brain tissue segmentation, we analyze the evolution of top-performing algorithms over time. Specifically, we compare the performance of the highest-ranked teams from the FeTA challenges in 2021, 2022, and 2024, evaluating segmentation accuracy across the dataset splits available in each respective year.
To extend the longitudinal comparison, we perform a retrospective evaluation of the 2022 winning method on the KCL dataset, first introduced as a test set in 2024. This is enabled by the 2022 winning team’s release of their Docker container [^9] https://hub.docker.com/r/fetachallenge22/feta-imperial-tum-2022-nnunet, allowing us to assess the generalization of a previously state-of-the-art solution to new, unseen data, and to identify both progress and persistent limitations.
2.7.2 Quantifying domain shifts
Domain shifts remain a key obstacle in fetal brain MRI analysis, often undermining model generalizability. These shifts arise from variations in subject demographics, imaging protocols, scanner types, and reconstruction methods (Dockès et al., 2021). In fetal imaging, GA notably affects brain morphology and contrast, while pathologies such as ventriculomegaly, for example, can significantly alter anatomical structure. Beyond biological and acquisition-related variability, low contrast or motion artifacts can degrade reconstruction quality, adversely affecting segmentation and biometry.
Is image quality a domain in itself? To assess whether image quality impacts model generalization, we manually rated the quality of all 180 test volumes using the protocol from Sanchez et al. (2024a) and explored the interaction of data quality with the performance of the submitted algorithms across the test data.
Comparing the impact of domain shift factors
To assess how domain shifts influence segmentation performance, we examine six key sources of variability: image quality, GA, condition (neurotypical or pathological), acquisition site, testing domain (seen vs. unseen during training), and the SRR method. These factors are summarized in Figure 3. To evaluate the influence of domain shift factors on segmentation performance, we trained a random forest regressor for each metric of interest (Dice, HD95, Volume Similarity, Euler Difference), using six dataset-level variables as input features. Target values were defined as the average metric scores across the top 3 teams. To estimate feature importance, we applied SHapley Additive exPlanations (SHAP) (Lundberg and Lee, 2017), which quantify the contribution of each feature by computing its average marginal effect across all possible feature combinations. This approach provides a unified and interpretable measure of how each factor affects performance.
Figure 3: Illustration of the sources of domain shifts in fetal brain MRI datasets of FeTA 2024. Demonstrated across gestational age (18 vs. 35 weeks), data quality (0.9 vs. 3.64), clinical condition (neurotypical vs. pathological), acquisition site (KISPI vs. CHUV), and SRR methods (IRTK vs. NiftyMIC). In each comparison, only the indicated domain is varied, while all other domains remain constant. Additional domains within each source, not shown here, are represented by circles.
3 Results
3.1 Algorithm description
We received 176 access requests for the KISPI training cohort hosted on Synapse during the challenge active period (May–July 2024). However, not all of these requests were related to the FeTA challenge, as the dataset is also available for broader research purposes. For the Vienna dataset, 53 data access applications were submitted, but only 30 applicants completed the data transfer agreement process and successfully received the data.
For the segmentation task, we received 16 valid submissions, all evaluated on the full test set. One team declined participation in this paper and was excluded from the analysis; results from the remaining 15 teams are presented. For the biometry task, 7 teams submitted results. One team (falcons) failed to generate valid outputs for all test cases and was penalized accordingly, as described in the Section 2.5. Notably, all biometry participants also submitted segmentation entries, leveraging segmentation outputs either as a preprocessing step or direct input for biometry estimation. A detailed description of each algorithm is provided in the supplementary materials (appendix A2) and summarized in Tables 2 and 3.
3.1.1 Common data and model augmentation strategies
Participants adopted a variety of approaches, with the majority utilizing 3D architectures—14 out of 16 for the segmentation task and 6 out of 7 for the biometry task. Across both tasks, two strategies were commonly used: data augmentation and model ensembling.
Data augmentation was universally applied, with all segmentation (16 teams) and biometry (7 teams) models using it. Standard transformations like flipping, rotation, scaling, and intensity shifts were common, while advanced methods, such as SynthSeg (Billot et al., 2023) or global intensity non-linear augmentations (GIN) (Ouyang et al., 2022), were used by 3 teams. Some teams also simulated domain-specific artifacts, including fetal motion and bias field.
Ensembling was a key approach in segmentation, used by 14 out of 16 teams. This included combining models trained on different cross-validation splits (4 teams) or using varied architectures, training setups, data orientations, or augmentation schemes (8 teams). Some also integrated pre- or post-processing models, like denoising autoencoders or skull-stripping (2 teams). Ensembling was less common in the biometry task, with only 2 out of 7 teams employing it, as most models built biometry predictions in a single pipeline on top of segmentation outputs.
3.1.2 Segmentation models
Among the 16 submissions, the most common architectures were nnU-Net Isensee et al. (2018) (9 teams) and U-Net (Ronneberger et al., 2015; Çiçek et al., 2016) (6 teams), often used as baselines. Many teams enhanced these models with attention mechanisms (Vaswani, 2017), residual connections (He et al., 2016), or ensembling. Others explored alternatives such as Swin Transformers (Liu et al., 2021), custom U-Net variants, or hybrid CNN–Transformer designs. Most models were developed in PyTorch (12 teams), with parameter counts ranging from 5M to 140M (median: 31M, mean: 44.8M).
Use of external data was limited to 5 teams, primarily leveraging dHCP data (Hughes et al., 2016; Edwards et al., 2022), fetal brain atlases (Gholipour et al., 2017; Uus et al., 2023a), or foundation models pretrained on large-scale image datasets (Roy et al., 2023).
3.1.3 Biometry models
All biometry models leveraged segmentation outputs, either as pre-processing, auxiliary, or core input. Two teams employed nnU-Net or U-Net variants for direct regression, while others used custom CNNs (1/7) or more complex architectures integrating attention mechanisms or hybrid designs (4/7). Prediction strategies varied across teams: two teams directly regressed biometry values; three teams predicted 3D landmark coordinates; and two teams generated 3D landmark heatmaps. In the latter two approaches, biometry values were subsequently computed using scripts provided by the organizers. Most teams used 3D models (6/7), implemented primarily in PyTorch (6/7), with one using TensorFlow. Three teams leveraged external data, such as dHCP and fetal brain atlases, or employed foundation models pre-trained on large-scale datasets.
Table 2: Summary of the algorithms submitted for the fetal brain tissue segmentation task.
Team name Model Architecture Deep Learning Framework Dim Data Augmentation Cross-Validation External Data Ensembling Original Aspects
cemrg Hybrid Cross Attention Swin Transformer and CNN PyTorch, nnUNet 3D Horizontal Flipping, Vertical flipping, scaling, normalization 5-fold No No The Cross Attention Transformer (CAT) block design.
CeSNE-DiGAIR 3D UNet MONAI 3D Deformable (SyN) registrations between couples of neurotypicall and pathological scans from the preprocessed training dataset. Skull-stripping with BOUNTI Not specified No Use of models for post-processing Denoising autoencoder for segmentation accuracy enhancement.
falcons 2D Attention Gated U-Net TensorFlow 2D Rotation, width/height shift, vertical/horizontal flip, zooming, brightness, gaussian noise, gaussian blurring Not specified 70 images from dHCP) Models with different architecture and (or) training data Series of preprocessing steps including brain extraction, alignment, and non-uniform intensity correction. Ensembling of models trained on different orientations (axial, sagital, coronal)
feta_sigma UxLSTMEnc, UNet nnUNet 3D Rotation, Scaling, Translation, Gaussian Noise, Mirror Transform. 5-fold No Models with different architecture and (or) training data Use of UxLSTM and ensembling with nnUnet, Background masking.
hilab nnU-Net PyTorch, nnUNet 3D Default nnU-Net augmentations, histogram equalization, differentiated probabilities for sample selection in random copy-paste augmentations, replication of challenging cases. 5-fold No Models with different architecture and (or) training data Applying histogram equalisation to 3D images, differentiated probabilities for sample selection in random copy-paste augmentations, strategically replicating challenging cases in the training data. Enesemble of 5 models with different hyperparameters and pre-processing settings.
jwcrad Residual-USE-Net PyTorch, MONAI 3D Rotation, scaling, translation, intensity shift, low resolution simulation. 5-fold No Model trained on differnt CV splits Custom auxiliary loss function based on transformation consistency.
LIT Attention UNet, nnUNet ResidualEncoderUNet PyTorch, nnUNet 3D Rotation, Scaling, Gaussian Noise, Gaussian Blur, Brightness Alteration, Contrast Adjustment, Low Resolution Simulation, Gamma Adjustment, Mirroring 6-fold No Model trained on differnt CV splits Custom brain mask detection with Attention Unet
lmrcmc nnUNet, SegVol nnUNet, MONAI 3D nnUNet: default, SegVol: flip, ScaleIntensity, ShiftIntensity, GibbsNoise, BiasField, KSpaceSpikeNoise and Affine augmentation; with SLAug. Not specified No Models with different architecture and (or) training data Ensemble of U-Net and a foundation model, use of the SegVol model in fetal brain segmentation.
mic-dkfz U-Net (nnUNet), U-Net with Residual encoder nnUNet 3D randomized; blur, gaussian noise, spatial (rotation, scaling, flipping), brightness, contrast, low-resolution simulation, gamma, sharpening, blank rectangle 5-fold Yes (pre-training in MultiTalent) Models with different architecture and (or) training data
Pretraining with MultiTalent on a
collection of publicly available datasets.
Ensemble of 3 nn-Unet configurations
with different data augmentations .
paramahir_2023 3D UNet (segmentation), custom UNet-based (biometry) MONAI 3D Random Flipping, Random Rotation, Random Intensity Shifts Not specified No No Combination of segmentation and biometry prediction in a unified pipeline.
pasteurdbc MedNeXt_L and nnUNet nnUNet 3D RandomScaling, RandomRoatation, RandomAdjustContrast, RandFlip 5-fold Multi-modal multi-organ medical image datasets used in the pre-trained MedNeXt_L foundational model Models with different architecture and (or) training data Used additional datasets with CT and brain MRI images for model pre-training
qd_neuroincyte Swin UNETR MONAI 3D Random sliding window, flipping, 1% gaussian noise, rigid rotation of ± 25° around all axes, random shifting ± 5 mm along all axes. Not specified No Use of models for post-processing Brain masking for vienna
unipd-sum-aug 2D Swin-UMamba PyTorch, nnUNetv2, Monai 2D TorchIO transforms and GIN techniques, pair-wise co-registration, affine and rigid transforms using the Advanced Nomalization Tools. 5-fold Model pre-trained on ImageNet Model trained on differnt CV splits Pretrained on imageNet repository and using GIN.
UPFetal24 nnU-Net ResEncL nnUNetv2 3D Default nnU-Net augmentations; differentiated by specific data augmentations for each of the three models. 5-fold for config dHCP and fetal atlasses Models with different architecture and (or) training data Data augmentation strategies and ensembling of models
ViCOROB nnUNet nnUNet, PyTorch 3D Random bias field, motion artifacts, low-resolution simulation, SynthSeg-inspired T2w image synthesizer 3-fold No Model trained on differnt CV splits SynthSeg-inspired T2w image synthesizer, Sharpness-Aware Minimization (SAM) optimizer
Table 3: Summary of the algorithms submitted for the biometry estimation task.
Team name Architecture Dimensionality Original Aspects External datasets Framework/languange
qd_neuroincyte SwinUnetr 3D Relies on segmentation. Predict landmark heat maps only using the segmentation maps and then calculate biometry. No additional data was used Pytorch 2.2.2
CeSNE-DiGAIR CNN 3D Relies on segmentation. Predict the keypoints given the segmentation. No additional data was used PyTorch Version 2.4.0
jwcrad Residual-USE-Net 3D Relies on segmentation. Uses the segmentation maps to localize and preprocess the input images by masking and cropping the original 3D image. Predict landmark heat maps using the preprocessed images and then calculate biometry. No additional data was used PyTorch 2.2.2
pasteurdbc MedNeXt_L nnUNet 3D Use of a pre-trained foundational model. Yes (for the pre-trained MedNeXt_L foundational model, multi-modal multi-organ medical image datasets)
falcons Attention Gated U-Net 2D Relies on segmentation. Predict the biometry values directly Yes (+70 images from dHCP) Tensorflow(2.10.0) FMRIB Software Library(FSL 6.0), CIVET(2.1.0), Advanced Normalization Tools(ANTs), Scikit-learn (1.5.1)
feta_sigma nnUNet, UxLSTMEnc 3D Ensemble network of nnUnet and UxLSTMEnc. No additional data was used PyTorch
paramahir_2023 UNet 3D Relies on segmentation. Predict the biometry values by regressing the U-Net features. No additional data was used PyTorch 2.3 -
3.2 FeTA 2024 results
3.2.1 Brain tissue segmentation ranking
Segmentation performance overview
Figure 4 highlights performance across sites and metrics, revealing a general performance plateau among top methods. For most teams, average Dice scores stabilized around 0.8-0.82, HD95 around 2.8-2.1, and VS around 0.9-0.92, while the ED showed wider variability (ranging from 20 to 40), highlighting its sensitivity to topological inaccuracies.
Site-specific trends
Despite being introduced in this edition as a new low-field, out-of-domain dataset, KCL showed the best segmentation performance. In contrast, KISPI yielded the lowest performance, even though it was part of the previous editions’ training and testing data. Across metrics, UCSF and KISPI displayed higher interquartile ranges, particularly for Dice, HD95, and VS, reflecting greater variability across methods. Some teams (falcons, qd_neuroincyte) experienced performance drops on sites that use NiftyMIC SRR, like UCSF or VIEN, with Dice scores dropping to 0.38–0.44 compared to 0.76–0.83 on other sites.
Label-specific trends
SGM, GM, and BS were consistently the most challenging labels to segment across all teams, as shown by lower performance metrics in Supplementary Materials A4. Among the top three models, Dice scores dropped from an average (across all labels) of 0.82 to 0.80 for SGM, 0.79 for BS, and 0.74 for GM. HD95 increased from 2.24 to 3.6 for BS and 3.0 for SGM, while VS declined from 0.92 to 0.86 for SGM and 0.88 for BS. GM also showed a marked increase in ED, from 33.14 to 137, reflecting a significant loss in topological accuracy.
Figure 4: Segmentation performance by site and evaluation metric. In each subplot, teams are ranked from left to right based on their average performance across all labels for the given metric (best to worst). Team colors are consistent across plots and correspond to the legend.
Ranking summary
Table 4 presents the aggregated average metrics and rankings per team  [^10] The rankings originally published on the website and announced during the MICCAI challenge differ from those presented in this paper due to a change in the ED estimation method. Specifically, we updated the way the ground-truth Euler characteristic is determined. In the original rankings, it was computed based on the manual segmentations. In the current results, it is calculated from manually defined topological properties (see 2.4). Manual segmentations were created via interpolation and because many structures were not segmented on every slice, the resulting ground-truth segmentations contained numerous topological errors (e.g., holes, disconnected components). As a result, they did not reliably represent the expected topological properties of the anatomical structures. To address this, we now use manually specified topological values to calculate the ground-truth EC.. Qualitative examples of the segmentations are provided in Supplementary Appendix A9. Notable rank discrepancies across metrics highlight their complementary nature. Figure 5 provides a more granular view, showing single-metric rankings across different sites and tissue labels. Dice score rankings remained relatively consistent across submissions and anatomical regions, while ED rankings showed greater variability, both across tissues and sites, reinforcing the importance of using multiple metrics to capture distinct aspects of segmentation quality.
Figure 5: Detailed submissions rankings. Top: Across labels. Bottom: Across sites (circles indicated in-domain splits and crosses out-of-domain splits). Teams are sorted in each subplot by their overall ranking in the segmentation task, from the best to the worst.
Figure 6 further illustrates the added value of topological metrics. In a comparative example, mic-dkfz-feta24 achieves similar Dice and lower HD95 scores but poorer ED and VS, suggesting that voxel-level agreement alone may not suffice for tasks requiring topologically accurate surfaces, such as morphological analysis.
Figure 6: Segmentation results and reconstructed cortical GM surfaces for two representative fetal cases from cesne-digair and mic-dkfz-feta24, visualized using ITK-SNAP (Yushkevich et al., 2006). The top row (fetus at 22 weeks GA) illustrates that higher Dice scores sometimes correspond to smaller topological errors. However, the bottom row (fetus at 30 weeks GA) demonstrates significant topological issues (e.g., holes, fragmentation) in the mic-dkfz-feta24 surface, despite comparable Dice and HD95 values. This underscores the need for additional topological and structural metrics, such as ED, to comprehensively evaluate segmentation quality, as metrics like Dice or HD95 alone are insufficient to capture topological accuracy.
Per-tissue and condition analysis
Extended performance results split by site, tissue label, and pathology status are available in supplementary materials (sections A3, A4, and A5, respectively).
Table 4: Segmentation ranking and average metrics
Team Dice HD95 VS ED Mean rank Final rank
Rank Value Rank Value Rank Value Rank Value
cesne-digair 8 0.816 3 2.317 1 0.929 1 20.921 3.25 1
mic-dkfz-feta24 1 0.828 2 2.224 3 0.918 8 37.206 3.50 2
vicorob 2 0.825 1 2.187 2 0.920 11 41.293 4.00 3
feta_sigma 3 0.822 7 2.430 5 0.914 4 31.710 4.75 4
cemrg_feta 4 0.822 10 2.836 4 0.916 7 34.382 6.25 5
upfetal24 5 0.820 6 2.412 6 0.913 9 39.967 6.50 6
hilab 7 0.816 8 2.434 9 0.911 3 30.123 6.75 7
lit 10 0.808 5 2.391 8 0.911 10 40.085 8.25 8
lmrcmc 11 0.805 11 3.179 7 0.913 5 32.751 8.50 9
unipd-sum-aug 9 0.811 4 2.332 10 0.909 13 46.668 9.00 10
pasteurdbc 6 0.817 9 2.474 11 0.909 12 41.521 9.50 11
jwcrad 12 0.769 12 3.569 12 0.886 2 29.744 9.50 11
qd_neuroincyte 13 0.681 13 10.441 13 0.827 6 34.295 11.25 13
falcons 14 0.628 14 11.040 14 0.765 14 100.729 14.00 14
paramahir_2023 15 0.040 15 80.757 15 0.337 15 1416.515 15.00 15
3.2.2 Biometry ranking
Performance across sites and measurement
Figure 7 summarizes model performance per site and biometric measurement, with detailed values available in the supplementary materials section A6.
Figure 7: Biometry results per site (top) and label (bottom) for all teams participating in the Task 2 together with the GA baseline model ([GA]) and the inter-rater variability [inter-rater]. Teams are sorted in ascending order for each subplot independently, based on their mean MAPE for a given site or label.
VIEN was the most challenging site, where no method outperformed the [GA] baseline (MAPE: 0.106$\pm$0.112), including the best-performing teams cesne-digair and jwcrad, which reached similar error levels. In contrast, KISPI emerged as the least challenging, with all three top teams exceeding the baseline. Across KCL, UCSF, and CHUV, only two teams per site (out of the top 3: jwcrad, feta_sigma, cesne_digair) achieved better-than-baseline performance. Measurement-wise, LCC, HV, and TCD were consistently more difficult, with HV and LCC showing the highest MAPE across all teams and raters. In contrast, sBIP and bBIP were among the best estimated. Notably, only jwcrad surpassed the baseline across all measurements, while a few others, including feta_sigma and pasteurdbc, did so on selected metrics.
Although multiple teams performed comparably on individual metrics, the clear winner in the ranking (see Table 5) was jwcrad, demonstrating consistent superiority across both site and measurement variations.
Table 5: Metrics and ranking for the biometry estimation task sorted by the final MAPE. [GA] and [inter-rater] entries do not represent participating models, thus their rank is marked as *
Team LCC HV bBIP sBIP TCD Final MAPE Final rank
MAPE Rank MAPE Rank MAPE Rank MAPE Rank MAPE Rank
[inter-rater] 9.59 * 8.04 * 3.28 * 1.49 * 4.89 * 5.38 *
jwcrad 11.15 1 10.32 2 5.43 2 4.78 3 7.21 2 7.72 1
[GA] 12.75 3 11.26 3 6.82 5 6.47 5 10.80 3 9.56 *
cesne-digair 17.72 4 9.82 1 4.02 1 4.74 2 12.34 4 9.59 2
feta_sigma 12.59 2 11.55 4 5.74 3 5.54 4 13.66 5 9.76 3
pasteurdbc 20.47 5 43.48 7 6.51 4 3.74 1 5.43 1 15.83 4
paramahir_2023 28.48 6 29.35 5 26.13 7 25.46 6 30.78 6 28.03 5
falcons 34.88 8 46.25 8 24.62 6 28.13 7 36.72 7 34.09 6
qd_neuroincyte 32.78 7 42.84 6 38.41 8 37.83 8 47.92 8 40.07 7
Robustness in pathological vs. neurotypical condition
To assess model generalizability, we compared biometry performance between neurotypical and pathological brains (see Figure 8). While most measurements did not reveal statistically significant differences between groups, bBIP showed better accuracy in the healthy cohort, particularly at VIEN. Conversely, UCSF results suggested slightly better performance for pathological subjects.
Figure 8: Biometry results for healthy and pathological subjects across labels and sites for the winning team jwcrad. Asterisks above the boxplot indicate statistically significant differences between the two groups (p $<$0.05, Mann-Whitney test).
In summary, the best-performing method, jwcrad, came within 9% of expert agreement for some measurements (e.g., TCD). However, for others like bBIP, its results differed from the expert range by as much as 60%. This reveals important gaps where automated biometry methods still fall short, especially in pathological cases.
3.3 Segmentation performance across challenge editions (2021, 2022 and 2024)
Over the three editions of the FeTA challenge, the segmentation task has expanded both in terms of dataset size (from 40 to 180 test cases) and site diversity (from 1 to 5 imaging centers). To evaluate progress over time, we compared segmentation performance across the years 2021, 2022, and 2024, focusing on common testing sites. Table 6 summarizes aggregated metrics for the top-performing teams each year (cesne-digair for 2024, FIT_1 for 2022 and NVAUTO for 2021), and Figure 9 provides a visual overview of mean scores per label and site across years, with markers for statistically significant differences.
Table 6: Mean and standard deviation (mean±$\pm$±std) for different metrics across years and splits over all labels. Bold and * highlight the years that have statistically significant improvement in the values compared to the previous year. ED was not estimated in FeTA 2021
Year Site Dice HD95 VS ED
2021 KISPI 0.79$\pm$0.16 2.81$\pm$3.43 0.89$\pm$0.16 not estimated
2022 CHUV 0.81$\pm$0.09 2.33$\pm$1.68 0.88$\pm$0.10 77.43$\pm$168.91
KCL 0.87$\pm$0.05 1.46$\pm$0.57 0.95$\pm$0.05 28.61$\pm$53.89
KISPI 0.77$\pm$0.18 3.17$\pm$4.16 0.87$\pm$0.18 18.98$\pm$54.56
UCSF 0.84$\pm$0.06 2.02$\pm$1.44 0.95$\pm$0.05 18.73$\pm$44.52
VIEN 0.84$\pm$0.08 1.87$\pm$1.46 0.95$\pm$0.06 32.20$\pm$67.05
2024 CHUV 0.83$\pm$0.06 2.23$\pm$1.40 0.93$\pm$0.06* 29.10$\pm$51.56*
KCL 0.86$\pm$0.05 1.69$\pm$0.52 0.95$\pm$0.04 6.26$\pm$13.21*
KISPI 0.78$\pm$0.15 2.95$\pm$2.86 0.89$\pm$0.14 9.21$\pm$17.84
UCSF 0.82$\pm$0.07 2.13$\pm$1.31 0.94$\pm$0.05 14.57$\pm$25.90
VIEN 0.81$\pm$0.09 2.27$\pm$1.69 0.95$\pm$0.05 38.13$\pm$93.48
KISPI split (2021–2024). This is the only site included in all three editions. No statistically significant improvement over the years was observed across the tracked metrics: Dice (0.79$\pm$0.16 $\rightarrow$ 0.77$\pm$0.18 $\rightarrow$ 0.78$\pm$0.15), HD95 (2.81$\pm$3.43 $\rightarrow$ 3.17$\pm$4.16 $\rightarrow$ 2.95$\pm$2.86), and VS (0.89$\pm$0.16 $\rightarrow$ 0.87$\pm$0.18 $\rightarrow$ 0.89$\pm$0.14). The only statistically significant change occurred in the VS metric for the GM label between 2021 and 2022 (0.96 $\rightarrow$ 0.94), but no consistent improvement was found in subsequent years or for other labels.
Other sites (2022–2024). For sites such as CHUV, UCSF, and VIEN—which were included in both 2022 and 2024—no consistent improvement was observed across metrics or tissue labels. While some metrics showed statistically significant changes, these were isolated and not consistent across sites, making it difficult to interpret them as evidence of overall progress in segmentation performance. Notably, ED improved substantially for CHUV and KCL. However, this improvement may be partially influenced by the inclusion of ED in the 2024 ranking, which favored algorithms with better topological performance. At CHUV, both VS and ED were significantly better than in 2022, but the other two metrics did not show similar trends.
Overall, although methods have become more sophisticated and the data more diverse, performance has not consistently improved across editions.
Figure 9: Segmentation performance improvement over the three editions of the FeTA Challenge. The superscript number above each label indicates whether the performance for a particular year and site-label-metric combination was statistically significantly better compared to all other available years. A superscript ”2” indicates that the results for 2022 were the best, while a superscript ”4” indicates that the results for 2024 were the best.
3.4 Domain shifts evaluation
3.4.1 Impact of image quality on performance
The impact of image quality on model performance as determined by computing the conditional mean across quality ratings ($\mathbb{E}[f(x)|\text{Quality}]$) is shown in the right-most column in Figure 10. We see a clear effect of image quality on Dice, with a generally increasing Dice with the increasing image quality, amounting to a change from 0.75 Dice on average for the lowest quality data (with scores close to 1) and an average quality close to 0.85 for the highest quality data. Results using HD95 and VS generally align with the ones from Dice, except for GA and quality. The relationship is, however, not as clear for ED, although best quality images tend to yield the smallest ED.
A more detailed analysis of the correlation between quality and the difference scores in the supplementary materials A7 showed a generally high Pearson correlation between quality and Dice ($r$ = 0.5-0.7) for all sites except KISPI-mial ($r$=0.4) and USCF-nmic ($r$=0.06 – no correlation). The same trends, although weaker, were observed for HD95 and VS, except CHUV-mial and HD95, which had virtually no correlation ($r$=0.05). Results for ED showed no clear pattern, and larger correlations ($r$=0.3-0.4) were not statistically significant.
Figure 10: Conditional mean plots for Dice, HD95, VS and ED, for different domain shift factors: pathology, site and SR, GA and image quality. The conditional mean shows how a given metric deviates from the global expected performance ($\mathbb{E}[f(x)]$) when a specific variable is used for conditioning.
Figure 11: SHAP value distribution across the data for segmentation metrics. Compared to conditional means (Fig. 10), SHAP values are the attribution of the impact on Dice or ED of different factors. Blue dots correspond to lower values of a variable (i.e. low GA, low quality), and red ones to higher values.
3.4.2 Relative contribution of domain-shift sources
Figure 10 displays the conditional means across different factors. The analysis revealed a pronounced site-SR effect: for example, the KISPI-mial site produced notably lower Dice scores, whereas the CHUV-mial site was associated with higher ED values. In addition, gestational age (GA) significantly affected both Dice and ED scores. Similar trends can be found in the supplementary materials (appendix A7), although HD95 scores appear to be less influenced by GA.
Figure 11 presents a SHAP analysis for all metrics using only image-level descriptors—namely, image quality, subject condition, and gestational age. (We excluded Site-SR from this analysis because its dependence on the other variables could lead to misleading SHAP values under the assumption of feature independence (Mase et al., 2019).) Overall, the SHAP analysis summarizes how these factors influence the Dice and ED scores: image quality generally has the largest impact, followed by GA (except for HD95). Although the pathological status of a subject generally has a small effect, we observed that severely pathological cases often have lower GA, which might introduce confounding. The plot’s color coding further confirms that higher image quality and GA are associated with increased Dice scores—for example, poor quality data may result in about –0.1 Dice, compared to an average of +0.05 Dice for good quality data.
4 Discussion
4.1 FeTA 2024 results and ranking
The multi-site, multi-task design of this challenge offered a unique opportunity to evaluate the progress and robustness of fetal brain image analysis algorithms. We summarize the main observations below.
Submitted methods overview
Analyzing solutions of the best 3 teams for the segmentation task, we see that all of them used 3D models, specifically U-Nets or nnU-Nets. External data did not play a major role, with two of the top three teams relying solely on the provided dataset. Notably, the first-place team, cesne-digair, uniquely incorporated a denoising autoencoder (Larrazabal et al., 2020) to enhance segmentation accuracy. This approach significantly boosted their performance, leading to a 50% improvement in the ED metric compared to the second-best team. This gain was particularly important for topological correctness, as no other metric showed such a large performance gap. All top-three teams applied extensive data augmentation, including techniques like SynthSeg (Billot et al., 2023), combinations of standard augmentations, and deformable registration between pathological and healthy subjects to simulate greater anatomical diversity, as well as model ensembling.
The top three biometry teams each approached estimation differently—using keypoint regression, landmark heatmap segmentation, and other methods—but all relied on the results of the tissues segmentation and 3D models. Their model architectures varied (CNN, U-Net, and Transformer), indicating that no standard baseline has yet emerged for this task. The diversity in methods suggests that the field is still exploring the most effective strategies for biometry estimation.
Topology metric is a valuable add-on
In the brain tissue segmentation task, the introduction of the topology-aware metric provided meaningful complementary insights beyond traditional overlap-based measures. Despite the architectural diversity and growing methodological sophistication of the submitted approaches, the performance differences among the top teams were minimal, with Dice scores showing tight clustering, suggesting that gains in segmentation accuracy may be reaching a plateau. Differences in team rankings across evaluation metrics (Dice, HD95, VS, ED) highlight the need to consider complementary metrics beyond voxel-wise overlap. Introducing ED as a ranking metric provided a more nuanced assessment of the segmentation quality. This is reflected in Table 6, where we see a marked improvement in ED. While teams did not specifically optimize their models for topological consistency, the new ranking scheme allowed us to discriminate between methods that otherwise had very similar performances (Table 4). Evaluation noise, where performance variations across testing sets is larger than the difference across top-performing methods, is a well-known problem in medical imaging challenges (Varoquaux and Cheplygina, 2022), and the introduction of an additional ranking metric allowed for selection methods with desirable properties. Further validation on clinical tasks leveraging surface extraction(Clouchoux et al., 2011; Yehuda et al., 2023) would be needed to truly see the potential of encouraging topological consistency in the FeTA challenge.
Low-field MRI tissue segmentation quality is encouraging
The newly introduced 0.55T data from KCL provided an unexpected insight: it consistently achieved the highest segmentation accuracy across all sites. However, it is important to note that, to reduce domain shifts, we retrospectively selected high-quality reconstructions using a version of SVRTK specifically tailored for low-field MRI data(Uus et al., 2020, 2024). This careful case selection, paired with a very recent SR pipeline, might have positively biased the performance for this cohort. As such, performance on more challenging low-field cases remains to be fully assessed. Nevertheless, these results are encouraging for two reasons: progress in SR pipelines (Uus et al., 2024; Xu et al., 2023) means that more challenging cases will be successfully reconstructed, and that image quality will generally increase (Sanchez et al., 2024a; Uus et al., 2023b). Low-field MRI systems hold significant promise for expanding access to prenatal imaging, particularly in low- and middle-income countries. When combined with good image quality and advances in automatic fetal exam planning (Neves Silva et al., 2024), they could meaningfully enhance prenatal care in resource-limited settings.
Automated biometry needs strong baselines to ensure meaningful progress
To reflect current clinical practice and bridge the gap between routine 2D fetal brain assessments and emerging 3D imaging techniques, we introduced a new biometry task focused on 2D brain measurements—key clinical indicators traditionally used to assess fetal neurodevelopmental status (Tilea et al., 2009; Lamon et al., 2024). One of the striking results of this first edition is that most submissions did not manage to outperform a simple model, that predicts biometry solely based on gestational age, completely ignoring image information. Most teams built biometry estimators atop segmentation outputs, potentially propagating segmentation errors, particularly in smaller and more complex structures such as vermis and corpus callosum measurements. Importantly, the biometric measurements used in this challenge were derived from clinical 2D protocols. While these can be estimated from 3D volumes, they were originally designed for manual 2D evaluation. This suggests that alternative measurements, specifically tailored to leverage the full spatial context of 3D SRR images, may offer even more informative and robust indicators of fetal neurodevelopment, though such approaches remain largely unexplored. Nevertheless, this first competition confirms, once more, the need to have strong baseline models and validation procedures (Eisenmann et al., 2023), and that deep learning might not always be the optimal solution (Grinsztajn et al., 2022).
4.2 FeTA challenge in perspective
FeTA across the years
A retrospective analysis of FeTA challenge results over the years revealed no statistically significant improvements in performance metrics, with the exception of ED at two out of five sites. Similarly, no notable improvement was seen at the label level, with GM, SGM, and BS consistently remaining the most challenging structures to segment. GM is particularly difficult due to its very thin appearance in fetal brains, where partial volume effects and complex surface morphology make it especially prone to topological segmentation errors, leading to significantly higher ED values compared to other labels. Moreover, both GM and SGM have inherently low tissue contrast in MRI, making them harder to distinguish accurately (Prayer et al., 2006). These challenges are further illustrated in Supplementary Materials A9.1, which provide qualitative examples showing that most segmentation errors occur in regions corresponding to GM, SGM, and BS. This outcome is not entirely unexpected, as most top-performing teams relied on similar 3D architectures—primarily 3D U-Net (Çiçek et al., 2016) and nnU-Net (Isensee et al., 2018)—enhanced with extensive data augmentation and model ensembling. These findings suggest that incremental architectural modifications or model engineering alone are unlikely to yield substantial gains, aligning with trends observed in other challenges where U-Net-based approaches often outperform more complex alternatives (Eisenmann et al., 2023). While these techniques help mitigate certain domain shifts related to scanner differences or pathological variations, some cases remain persistently difficult across all methods. Addressing these harder cases may require deeper domain expertise and a shift toward a more data-centric approach, prioritizing data quality, annotation consistency, and dataset diversity as core components of model development (Sambasivan et al., 2021; Zha et al., 2023).
Sources of domain shifts
Domain shifts are widely recognized as a key challenge for deep learning methods in medical imaging (Dockès et al., 2021; Wiles et al., 2021; Richiardi et al., 2025), yet the specific sources of these shifts are rarely disentangled. In our analysis, though not causal, we observed that image quality had the strongest impact on generalization performance: moving from the lowest to the highest quality levels resulted in an average Dice score difference of approximately 0.10. In contrast, gestational age had a more modest effect, influencing Dice scores by about 0.05, while the scanning site contributed a difference of around 0.075 between the best- and worst-performing centers. Interestingly, pathology was the least influential factor, accounting for only about 0.008 in Dice variation. Additionally, because Dice is known to be biased toward larger structures (Maier-Hein et al., 2024), we also evaluated performance using a normalized Dice metric that accounts for label volume (Raina et al., 2023). As detailed in the supplementary materials section A8, the normalized Dice scores yielded rankings nearly identical to those based on standard Dice, indicating that structure size did not significantly distort the comparative performance of the reviewed algorithms, suggesting that while the size bias exists, its effect was uniform across methods.
Our results show that, despite pathological cases making up only about one-third of the training data, models were still able to generalize to pathological examples in the test set. While performance was slightly lower for pathological subjects in some datasets compared to healthy subjects, submitted models demonstrated the ability to correctly handle both healthy and pathological data. Given the rarity and wide variability of fetal pathologies (Attallah et al., 2019), expanding pathological datasets—whether through additional real cases or synthetic data (Dannecker et al., 2024; Kaandorp et al., 2025b)—will be crucial to narrowing this performance gap and improving overall model robustness, which is an important step toward real-world clinical deployment.
Overall, our findings suggest that technical and acquisition-related factors may play a more significant role in out-of-domain generalization than subject-level clinical variables. Still, further causal investigations (Castro et al., 2020) are needed to confirm these patterns and to avoid misinterpretation due to confounding factors.
4.3 Roadmap for future advancements in fetal brain MRI analysis
While many proposed solutions appear to be reaching a performance plateau, model-centric innovations still play an important role. That said, incorporating domain-specific augmentations and auxiliary learning objectives may lead to more impactful improvements than simply refining model architectures. For example, enforcing topological consistency within the loss function, as demonstrated by de Dumast et al. (2022); Li et al. (2023); Lux et al. (2024)—can help maintain anatomical plausibility in the predictions. Similarly, integrating uncertainty estimation provides a powerful way of identifying low-confidence predictions, which is particularly relevant in clinical decision-support systems. Several studies (Zenk et al., 2025; Molchanova et al., 2025) have demonstrated the utility of uncertainty-aware models for quality control in medical image segmentation.
Beyond model architecture, data-driven strategies hold substantial potential for improvement. A notable limitation of current solutions is the relatively modest use of external data, which has been largely limited to healthy subjects from datasets like dHCP or fetal brain atlases (Gholipour et al., 2017; Uus et al., 2023a; Price et al., 2019). Leveraging broader, more diverse datasets — especially those capturing rare or pathological conditions — could support more robust and clinically useful models, though curating and annotating pathological datasets is a huge endeavor.
Manually segmenting the fetal brain is a time-consuming and tedious task, susceptible to inter-rater variability (Payette et al., 2021), and the FeTA challenge data are not exempt from this issue. When comparing model performance to inter-rater variability, we observe that top-performing teams—achieving Dice scores around 0.82, HD95 around 2.2, and VS around 0.92—are approaching the best observed human agreement levels, previously estimated on a subset of data as $0.73\pm 0.15$ for Dice, $3.45\pm 2.34$ for HD95, and $0.86\pm 0.10$ for VS (Payette et al., 2023). This raises the intriguing possibility that some predictions may be more faithful to the underlying anatomy than the ground truth annotations, potentially leading to penalization of high-performing models (Valabregue et al., 2024a, b).
A promising direction to address the limitations of data diversity and annotation availability is data synthesis (Zalevskyi et al., 2024), particularly for generating rare or pathological fetal brain appearances. Recent work (Dannecker et al., 2024; Liu et al., 2024; Kaandorp et al., 2025a) has highlighted the potential of synthetic data to augment training and improve sensitivity to abnormal anatomy. Moreover, the strong influence of image quality on generalization performance underscores the need for better modeling of artifacts specific to fetal brain SR pipelines (Sanchez et al., 2024a). These efforts, in combination with foundation models and domain adaptation techniques, offer exciting prospects for enhancing model generalization across scanners, domains, and populations, ultimately helping to mitigate model drift and support the development of more trustworthy AI systems.
5 Conclusion
The FeTA 2024 challenge provided a valuable opportunity to evaluate the progress made in fetal brain segmentation since previous editions and to expand the scope toward new, clinically relevant tasks such as biometry. Our additional validation using the Euler difference metric showed that some existing methods can already produce topologically consistent segmentations. However, achieving this consistency more reliably, particularly through improved segmentation losses, remains an open area for further development. Likewise, the successful application of models to low-field data, with surprisingly strong performance, highlights both the advancements in recent super-resolution methods and the models’ capacity to generalize across diverse imaging settings.
In the biometry task, this first edition offered key insights, particularly on the importance of providing simple baseline models to guide participants. It also led to the emergence of a promising approach for automated biometry prediction.
As the field of fetal brain MRI analysis continues to evolve, FeTA 2024 emphasizes the need not only for more powerful and innovative models but also for building reliable and generalizable tools that can support real-world clinical adoption.
Declaration of competing interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Part of the challenge data is available on Synapse, as described in the paper.
Acknowledgments
This research was funded by the Swiss National Science Foundation (182602, 215641, IZKSZ3_218590), ERA-NET Neuron MULTI-FACT project (SNSF 31NE30 203977), UKRI FLF (MR/T018119/1), the Adaptive Brain Circuits in Development and Learning Project, University Research Priority Program of the University of Zürich; by the Vontobel Foundation; by the Anna Müller Grocholski Foundation and the Prof. Max Cloetta Foundation and DFG Heisenberg funding (502024488); we acknowledge the Leenaards and Jeantet Foundations as well as CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by CHUV, UNIL, EPFL, UNIGE and HUG.
Diego Fajardo-Rojas would like to acknowledge funding from the EPSRC Centre for Doctoral Training in Smart Medical Imaging (EP/S022104/1).
Lyuyang Tong, Bo Du and Jingwen Jiang would like to acknowledge funding from the National Key Research and Development Program of China under Grants 2023YFC2705700, the National Natural Science Foundation of China under Grants 62306217 and 62225113, the Postdoctoral Fellowship Program of CPSF under Grant Number GZC20231987, the China Postdoctoral Science Foundation under Grant Number 2024T170686 and 2024M752471, the Major Program (JD) of Hubei Province (2023BAA017), the Innovative Research Group Project of Hubei Province under Grants 2024AFA017.
Gerard Martí-Juan is supported by the project PCI2021-122044-2A, funded by the project ERA-NET NEURON Cofund2, by MCIN/AEI/10.13039/501100011033/ and by the European Union NextGenerationEU/PRTR.
R. Hamadache holds an IFUdG PhD grant from the University of Girona. R. Hamadache and X. Lladó received support by the PID2023-146187OB-I00 from the Ministerio de Ciencia e Innovación, Spain.
M.O. Candela-Leal, A. Gondova, and S. You received support from the National Institute of Neurological Disorders and Stroke (R01NS114087) and National Institute of Biomedical Imaging and Bioengineering (R01EB031170) of the National Institutes of Health (NIH).
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used Grammarly and ChatGPT (GPT 4o) to assist with spell checking, grammar refinement, and language clarity improvements. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.